A mixture of 50.0 g sulfur and 1.00x102 g of molecular chlorine gas reacts completely to form disulfur dichloride and sulfur dichloride. Find the mass of disulfur dichloride formed and determine the limiting and remaining reagents?


1
Expert's answer
2021-11-19T07:37:03-0500

50*1*10^2=50*10^2=50*100=5000